# Solar flare forecasting with foundational transformer models across image, video, and time-series modalities

### v1

@@ Abstract
Over the past decades, flare forecasting has evolved through several methodological stages, progressing from empirical and statistical approaches to physics-based modeling and, more recently, to machine learning techniques. The advent of high–spatial and high–temporal resolution multi-wavelength observations from modern solar observatories such as NASA’s Solar Dynamics Observatory (SDO, Pesnell et al. 2012), ESA/NASA’s Solar Orbiter (Müller et al., 2020), NASA’s Parkes Solar Probe (PSP, Raouafi et al. 2023), ESA’s Proba-3 (Zhukov et al., 2025), the Advanced Space-based Solar Observatory (ASO-S, Gan et al. 2023), ISRO’s Aditya-L1 (Tripathi et al., 2022), and NOAA’s Geostationary Operational Environmental Satellites (GOES) has enabled the development of deep learning models capable of forecasting directly from images, sequences, videos, and multimodal data. These approaches minimize the need for manual feature engineering and can potentially uncover complex spatio-temporal patterns associated with solar activity.
In this context, the goal of this work is to explore how recent state-of-the-art transformer-based foundational architectures, designed for different data modalities (images, videos, time series), perform on the solar flare forecasting problem. To support future research and reproducibility within the community, we carried out our analysis on a publicly available dataset and released both the developed software and fine-tuned models as open resources. While transformer-based video models have only recently been applied to flare forecasting (Li et al., 2025), this study represents, to the best of our knowledge, one of the first explorations of foundational models for time-series in this domain.
@@ 2 Related work
Traditional solar flare forecasting has relied on handcrafted physical and empirical predictors, such as magnetic field features, historical flare activity, and statistical models. The recent availability of high-resolution solar imagery and magnetogram data (e.g., from SDO/HMI111Helioseismic and Magnetic Imager) has facilitated the adoption of deep learning techniques for this task. In this section, we review the main deep learning approaches proposed to date, organized by data modality, and summarize their forecasting performance in terms of the True Skill Statistic (TSS) and the Heidke Skill Score (HSS) metrics (see Section 4.2 and B).
@@ 2.3 Time series-based approaches
Building on this foundation, Abduallah (2023) introduced SolarFlareNet, a more sophisticated hybrid model that integrates 1D-CNN, LSTM, and transformer encoder blocks to capture both short- and long-term dependencies in the SHARP parameters. Trained on 108-minute sequences of nine magnetic parameters sampled at 12-minute cadence, the model forecasted flares at various intensity thresholds (C, M, M5.0) and horizons (24, 48, and 72 hours). Using the start of each sequence as the reference time for labeling,222This differs from our approach, in which the last time step defines the forecast reference. their best model achieved a mean TSS of 0.84 0.03 for 24-hour M+ forecasting one of the highest values reported for time-series methods to date.
@@
For completeness, we also note several recent studies that have focused on high-cadence time series of soft X-ray fluxes (XRS) recorded by the GOES. Unlike the previously discussed works, which frame flare prediction as a classification problem (i.e., predicting the flare class), these analyses approach the task as a multi-step regression problem, aiming to forecast the future evolution of the XRS flux itself or related observables over specific time horizons. For instance, Dei & Fuentes (2020) compared three neural architectures 1D CNN, LSTM, and N-BEATS for forecasting X-ray fluxes at lead times ranging from 1 to 5 hours, using historical windows of 1, 2, and 3 times the prediction interval. Among these, the N-BEATS model consistently achieved the lowest mean squared error (MSE0.001 for the largest prediction window), outperforming both baseline estimators and CNN/LSTM networks. More recently, van der Sande & Muñoz-Jaramillo (2025) evaluated a set of traditional machine learning and deep learning models (Linear, MLP, Random Forest, and CNN) to predict XRS fluxes at horizons up to 48 h. Their models used a diverse set of input sources, including magnetograms, EUV images at multiple wavelengths, and scalar magnetic features. Forecasting accuracy was found to improve with longer prediction windows, with their best CNN model reaching MSE 0.3 after 24 h. From the predicted flux values, they also derived binary classification metrics for M1-class events, obtaining TSS and HSS values of approximately 0.6 for forecasting windows beyond 24 h.
@@ 3.1 Image dataset
To mitigate this imbalance, we created additional downsampled datasets by randomly reducing the number of No-Flare and C-class images while retaining at least one image per AR (so ARs with more images were more aggressively downsampled). The target number of images for each downsampled class was set approximately equal to the combined number of M+ samples. The resulting datasets, denoted with a ’_ds’ suffix (e.g., train_ds), and their class abundances are reported in Table 1. More strongly balanced versions, labeled ’_ds_bal’ (e.g., train_ds_bal), were produced by further downsampling No-Flare and C-class data to achieve a nearly 1:1 ratio between positive and negative examples. Unless otherwise specified, only the downsampled datasets were used for training the image-based forecasting models.
@@ 3.3 Time series dataset
The time series dataset was constructed from historical solar irradiance measurements collected by the GOES333https://www.ncei.noaa.gov/products/goes-r-extreme-ultraviolet-xray-irradiance444https://www.ncei.noaa.gov/products/goes-1-15/space-weather-instruments, covering nearly three decades of solar activity from 1995 onward. Specifically, we used the science-quality Level 2 (L2) series of 1-minute averaged soft X-ray fluxes in the long-wavelength band (0.1–0.8 nm; XRS-B channel). For each GOES satellite (08-18), the science fluxes were normalized by their corresponding daily background flux level (reported by GOES), producing a flux-ratio time series (science/background). These normalized series were then segmented into non overlapping 12h and 24h intervals (respectively containing 720 and 1440 entries with 1-minute cadence), each representing a distinct observation sample for forecasting. Samples with more than 70% missing data were discarded, while shorter data gaps were filled by linear interpolation. Finally, a logarithmic transformation was applied to the flux ratios.
We enriched the time series dataset using historical flare event information from the NOAA SWPC Event Reports555ftp://ftp.swpc.noaa.gov/pub/warehouse/, which provide start and end times, flare class, and associated active region identifiers (when available). For each XRS flux-ratio series, we generated a corresponding binary sequence with the same temporal cadence (1-minute), where each time step was assigned a value of 1 if a flare of any class occurred within that minute, and 0 otherwise.
@@ 4.1 Forecasting models
Our forecasting framework is implemented using the transformers666https://github.com/huggingface/transformers and PyTorch777https://github.com/pytorch/pytorch libraries and supports three distinct data modalities: images, videos, and time series. As illustrated in Fig. 3, each modality is processed by a dedicated backbone model designed to extract domain-relevant representations from its respective input type:
@@ 4.1.2 Video model
For video-based forecasting we considered VideoMAE888Video Masked Autoencoder, https://github.com/MCG-NJU/VideoMAE model (Tong et al., 2022), a transformer-based architecture specifically designed for spatio-temporal representation learning in videos. VideoMAE extends the masked autoencoding paradigm from images to videos by introducing tube masking, where spatio-temporal cubes of a video are randomly masked at an extremely high ratio (90–95%). The model is trained in a self-supervised manner to reconstruct the missing patches from the remaining visible tokens. This strategy forces the encoder to capture high-level structures and motion dynamics rather than relying on pixel-level redundancy, making it particularly suitable for solar magnetogram sequences in which the temporal evolution of active regions is essential.
@@ 4.1.3 Time series model
For time-series forecasting we employ Moirai999Masked EncOder-based UnIveRsAl TIme Series Forecasting Transformer, https://github.com/SalesforceAIResearch/uni2ts, a recently proposed foundation model for universal time-series forecasting (Woo et al., 2024). Moirai is built on a Transformer encoder architecture but introduces several innovations that make it particularly effective for heterogeneous, long, and irregularly sampled sequences. Its design combines (i) multi patch-size projections, which enable efficient handling of both high- and low-frequency data, (ii) Any-variate Attention, which flattens multivariate inputs and encodes both time and variate axes so that the model can flexibly process any number of input channels, and (iii) mixture distribution heads, which allow probabilistic forecasting with flexible distributions beyond simple Gaussian assumptions.
@@ 4.2 Model training and evaluation
All training experiments were conducted on the CINECA LEONARDO Booster supercomputing infrastructure101010https://www.hpc.cineca.it/systems/hardware/leonardo/, using a single compute node equipped with 4 GPUs (NVIDIA A100, 64 GB) and 1 CPU (Intel Xeon Platinum 8358 CPUs, 2.60 GHz), with 512 GB of RAM. The following subsection outlines the details of the adopted training strategy.
@@
Because the solar flare datasets are highly imbalanced toward negative (non-flaring) examples, we implemented a custom Hugging Face (HF) trainer capable of rebalancing the training data within each batch.111111When the number of negative samples greatly exceeds that of positive samples and the batch size is small, there is a high probability that a batch will be dominated or entirely composed of negative examples. The trainer supports three alternative class-weighted loss functions to mitigate class imbalance: the class-weighted binary cross-entropy loss (), the class-weighted focal loss (), and the score-oriented loss (SOL; ) introduced by Marchetti et al. (2022) and Guastavino et al. (2023). A detailed mathematical description of these loss functions is provided in A.
@@
For each sample with true target (with denotes the occurrence of a flare), the model outputs a probability 121212Model outputs for trained binary forecasters are single logits, converted to probabilities through a sigmoid activation function. representing the likelihood of a flare occurring within the 24-hour forecasting window. The predicted class is obtained by thresholding this probability at (=0.5 by default), i.e., . From the pairs we computed the entries of the confusion matrix:
@@
Binary forecasting performance was evaluated using the following metrics:
| F1-score | ||||
Here, () denotes the recall (or true positive rate/sensitivity), is the precision, the true negative rate (specificity), and the false positive rate (probability of false alarm). is the True Skill Statistic (or Hanssen & Kuipers’ discriminant, H&KSS), is the Heidke Skill Score, and is the Matthews correlation coefficient. Additional details and discussion of these metrics are provided in B.
@@ 5 Results
Table 6 summarizes the forecasting metrics obtained on the test set for the best-performing models selected through validation across all data modalities. For the image and video models, the table reports the mean and standard deviation of the scores computed over independent runs, each trained and evaluated on the same original downsampled data split from Boucheron et al. (2023), labelled as "default". In addition, metrics averaged over alternative splits each randomly drawn from the original dataset as described in Section 3 are also reported and labelled as "alt.". For the time series model, metrics are likewise averaged over independent data splits. The best performance metrics for each model are shown in Fig. 4 as colored markers blue for SigLIP2, orange for VideoMAE, and red for Moirai2 indicating the mean scores with their corresponding standard deviations. The SigLIP2 image model achieved TSS=0.646, HSS=0.261, and MCC=0.340, confirming the ability of vision–language transformers to extract meaningful magnetic features from single magnetograms. The VideoMAE model, trained on 16-frame magnetogram sequences, yielded slightly lower scores (TSS=0.604, HSS=0.207, MCC=0.288), suggesting a limited ability to fully exploit temporal information and capture pre-flare evolution, even when varying cadence intervals (32, 72). Overall, both the image and video models achieve slightly higher metric values when evaluated on the original data split compared to the alternative samples. The Moirai2 time-series model achieved the best overall performance, with TSS=0.736, HSS=0.691, and MCC=0.694, outperforming both the image and video models, particularly for the HSS and MCC metrics.
@@ Acknowledgements
The software code used in this work is publicly available under the GNU General Public License v3.0131313https://www.gnu.org/licenses/gpl-3.0.html on the GitHub repository https://github.com/inaf-oact-ai/solar-flare-forecaster/.
@@ Appendix B Forecasting metrics
It ranges from (perfectly wrong) through (no skill, random/constant forecast) to (perfect). TSS is widely used in space-weather because it is relatively insensitive to class imbalance, while other metrics like the F1-score are more affected when the proportion of the minority class (rare event) changes.

### v2

@@ Abstract
Over the past decades, flare forecasting has evolved through several methodological stages, progressing from empirical and statistical approaches to physics-based modeling and, more recently, to machine learning techniques. The advent of high–spatial and high–temporal resolution multi-wavelength observations from modern solar observatories such as NASA’s Solar Dynamics Observatory (SDO, Pesnell et al. 2012), ESA/NASA’s Solar Orbiter (Müller et al., 2020), NASA’s Parker Solar Probe (PSP, Raouafi et al. 2023), ESA’s Proba-3 (Zhukov et al., 2025), the Advanced Space-based Solar Observatory (ASO-S, Gan et al. 2023), ISRO’s Aditya-L1 (Tripathi et al., 2022), and NOAA’s Geostationary Operational Environmental Satellites (GOES) has enabled the development of deep learning models capable of forecasting directly from images, sequences, videos, and multimodal data. These approaches minimize the need for manual feature engineering and can potentially uncover complex spatio-temporal patterns associated with solar activity.
In this context, the goal of this work is to explore how recent state-of-the-art transformer-based foundational architectures, designed for different data modalities (images, videos, time series), perform on the solar flare forecasting problem. To support future research and reproducibility within the community, we carried out our analysis on a publicly available dataset and released both the developed software and fine-tuned models as open resources 111https://github.com/inaf-oact-ai/solar-flare-forecaster/222https://huggingface.co/inaf-oact-ai. While transformer-based video models have only recently been applied to flare forecasting (Li et al., 2025), this study represents, to the best of our knowledge, one of the first explorations of foundational models for time-series in this domain.
@@ 2 Related work
Traditional solar flare forecasting has relied on handcrafted physical and empirical predictors, such as magnetic field features, historical flare activity, and statistical models. The recent availability of high-resolution solar imagery and magnetogram data (e.g., from SDO/HMI333Helioseismic and Magnetic Imager) has facilitated the adoption of deep learning techniques for this task. In this section, we review the main deep learning approaches proposed to date, organized by data modality, and summarize their forecasting performance in terms of the True Skill Statistic (TSS) and the Heidke Skill Score (HSS) metrics (see Section 4.2 and B).
@@ 2.1 Image-based approaches
Other studies, such as Legnaro et al. (2025), explored the classification of sunspot magnetic type in solar active-region cutouts using ResNet18 and DeiT (Data-Efficient Image Transformer) architectures, demonstrating that combining magnetogram and continuum images can significantly enhance model performance.
@@ 2.2 Spatio-temporal and video-based forecasting approaches
Sun et al. (2022) employed 3D CNNs to model 24-hour SHARP magnetogram sequences (15 frames at a 96-minute cadence) for C+ and M+ flare prediction within the next 24 hours. To address class imbalance, they adopted a focal loss function and systematically explored the impact of the focal parameter . Their best configuration, with , achieved TSS = 0.826 and HSS = 0.667 for M+ forecasting.
@@
Grim & Gradvohl (2024) further extended transformer-based approaches by applying Improved Multiscale Vision Transformers (MViTv2) to 24-hour SDO/HMI magnetogram sequences (16 frames at a 96-minute cadence) for 24- and 48-hour flare forecasting. Using a weighted loss and oversampling to mitigate class imbalance, their best M+ model achieved TSS = 0.70 0.04 and HSS = 0.25 0.01 for the 24-hour horizon.
@@ 2.3 Time series-based approaches
Building on this foundation, Abduallah (2023) introduced SolarFlareNet, a more sophisticated hybrid model that integrates 1D-CNN, LSTM, and transformer encoder blocks to capture both short- and long-term dependencies in the SHARP parameters. Trained on 108-minute sequences of nine magnetic parameters sampled at 12-minute cadence, the model forecasted flares at various intensity thresholds (C, M, M5.0) and horizons (24, 48, and 72 hours). Using the start of each sequence as the reference time for labeling,444This differs from our approach, in which the last time step defines the forecast reference. their best model achieved a mean TSS of 0.84 0.03 for 24-hour M+ forecasting one of the highest values reported for time-series methods to date.
@@
For completeness, we also note several recent studies that have focused on high-cadence time series of soft X-ray fluxes (XRS) recorded by the GOES. Unlike the previously discussed works, which frame flare prediction as a classification problem (i.e., predicting the flare class), these analyses approach the task as a multi-step regression problem, aiming to forecast the future evolution of the XRS flux itself or related observables over specific time horizons.
For instance, Dei & Fuentes (2020) compared three neural architectures 1D CNN, LSTM, and N-BEATS for forecasting X-ray fluxes at lead times ranging from 1 to 5 hours, using historical windows of 1, 2, and 3 times the prediction interval. Among these, the N-BEATS model consistently achieved the lowest mean squared error (MSE0.001 for the largest prediction window), outperforming both baseline estimators and CNN/LSTM networks.
More recently, van der Sande & Muñoz-Jaramillo (2025) evaluated a set of traditional machine learning and deep learning models (Linear, MLP, Random Forest, and CNN) to predict XRS fluxes at horizons up to 48 h. Their models used a diverse set of input sources, including magnetograms, EUV images at multiple wavelengths, and scalar magnetic features. Forecasting accuracy was found to improve with longer prediction windows, with their best CNN model reaching MSE 0.3 after 24 h. From the predicted flux values, they also derived binary classification metrics for M1-class events, obtaining TSS and HSS values of approximately 0.6 for forecasting windows beyond 24 h.
@@ 3.1 Image dataset
To mitigate this imbalance, we created additional downsampled datasets by randomly reducing the number of No-Flare and C-class images while retaining at least one image per AR (so ARs with more images were more aggressively downsampled). The target number of images for each downsampled class was set approximately equal to the combined number of M+ samples. The resulting datasets, denoted with a "_ds" suffix (e.g., train_ds), and their class abundances are reported in Table 1. More strongly balanced versions, labeled "_ds_bal" (e.g., train_ds_bal), were produced by further downsampling No-Flare and C-class data to achieve a nearly 1:1 ratio between positive and negative examples. Unless otherwise specified, only the downsampled datasets were used for training the image-based forecasting models.
@@ 3.3 Time series dataset
The time series dataset was constructed from historical solar irradiance measurements collected by the GOES555https://www.ncei.noaa.gov/products/goes-r-extreme-ultraviolet-xray-irradiance666https://www.ncei.noaa.gov/products/goes-1-15/space-weather-instruments, covering nearly three decades of solar activity from 1995 onward. Specifically, we used the science-quality Level 2 (L2) series of 1-minute averaged soft X-ray fluxes in the long-wavelength band (0.1–0.8 nm; XRS-B channel). For each GOES satellite (08-18), the science fluxes were normalized by their corresponding daily background flux level (reported by GOES), producing a flux-ratio time series (science/background). These normalized series were then segmented into non overlapping 12h and 24h intervals (respectively containing 720 and 1440 entries with 1-minute cadence), each representing a distinct observation sample for forecasting. Samples with more than 70% missing data were discarded, while shorter data gaps were filled by linear interpolation. Finally, a logarithmic transformation was applied to the flux ratios.
We enriched the time series dataset using historical flare event information from the NOAA SWPC Event Reports777ftp://ftp.swpc.noaa.gov/pub/warehouse/, which provide start and end times, flare class, and associated active region identifiers (when available). For each XRS flux-ratio series, we generated a corresponding binary sequence with the same temporal cadence (1-minute), where each time step was assigned a value of 1 if a flare of any class occurred within that minute, and 0 otherwise.
@@ 4.1 Forecasting models
Our forecasting framework is implemented using the transformers888https://github.com/huggingface/transformers and PyTorch999https://github.com/pytorch/pytorch libraries and supports three distinct data modalities: images, videos, and time series. As illustrated in Fig. 3, each modality is processed by a dedicated backbone model designed to extract domain-relevant representations from its respective input type:
@@ 4.1.2 Video model
For video-based forecasting we considered VideoMAE101010Video Masked Autoencoder, https://github.com/MCG-NJU/VideoMAE model (Tong et al., 2022), a transformer-based architecture specifically designed for spatio-temporal representation learning in videos. VideoMAE extends the masked autoencoding paradigm from images to videos by introducing tube masking, where spatio-temporal cubes of a video are randomly masked at an extremely high ratio (90–95%). The model is trained in a self-supervised manner to reconstruct the missing patches from the remaining visible tokens. This strategy forces the encoder to capture high-level structures and motion dynamics rather than relying on pixel-level redundancy, making it particularly suitable for solar magnetogram sequences in which the temporal evolution of active regions is essential.
@@ 4.1.3 Time series model
For time-series forecasting we employ Moirai111111Masked EncOder-based UnIveRsAl TIme Series Forecasting Transformer, https://github.com/SalesforceAIResearch/uni2ts, a recently proposed foundation model for universal time-series forecasting (Woo et al., 2024). Moirai is built on a Transformer encoder architecture but introduces several innovations that make it particularly effective for heterogeneous, long, and irregularly sampled sequences. Its design combines (i) multi patch-size projections, which enable efficient handling of both high- and low-frequency data, (ii) Any-variate Attention, which flattens multivariate inputs and encodes both time and variate axes so that the model can flexibly process any number of input channels, and (iii) mixture distribution heads, which allow probabilistic forecasting with flexible distributions beyond simple Gaussian assumptions.
@@ 4.2 Model training and evaluation
All training experiments were conducted on the CINECA LEONARDO Booster supercomputing infrastructure121212https://www.hpc.cineca.it/systems/hardware/leonardo/, using a single compute node equipped with 4 GPUs (NVIDIA A100, 64 GB) and 1 CPU (Intel Xeon Platinum 8358 CPUs, 2.60 GHz), with 512 GB of RAM. The following subsection outlines the details of the adopted training strategy.
@@
Because the solar flare datasets are highly imbalanced toward negative (non-flaring) examples, we implemented a custom Hugging Face (HF) trainer capable of rebalancing the training data within each batch.131313When the number of negative samples greatly exceeds that of positive samples and the batch size is small, there is a high probability that a batch will be dominated or entirely composed of negative examples. The trainer supports three alternative class-weighted loss functions to mitigate class imbalance: the class-weighted binary cross-entropy loss (), the class-weighted focal loss (), and the score-oriented loss (SOL; ) introduced by Marchetti et al. (2022) and Guastavino et al. (2023). A detailed mathematical description of these loss functions is provided in A.
@@
For each sample with true target (with denotes the occurrence of a flare), the model outputs a probability 141414Model outputs for trained binary forecasters are single logits, converted to probabilities through a sigmoid activation function. representing the likelihood of a flare occurring within the 24-hour forecasting window. The predicted class is obtained by thresholding this probability at (=0.5 by default), i.e., . From the pairs we computed the entries of the confusion matrix cin the table below:
@@
Binary forecasting performance was evaluated using the following metrics: True Skill Statistic (), Heidke Skill Score (), Matthews correlation coefficient (). Additional details and discussion of these metrics are provided in B.
@@ 5 Results
Table 6 summarizes the forecasting metrics obtained on the test set for the best-performing models selected through validation across all data modalities. For the image and video models, the table reports the mean and standard deviation of the scores computed over independent runs, each trained and evaluated on the same original downsampled data split from Boucheron et al. (2023), labelled as "default". In addition, metrics averaged over alternative splits each randomly drawn from the original dataset as described in Section 3 are also reported and labelled as "alt.". For the time series model, metrics are likewise averaged over independent data splits.
The best performance metrics for each model are shown in Fig. 4 as colored markers blue for SigLIP2, orange for VideoMAE, and red for Moirai2 indicating the mean scores with their corresponding standard deviations. The SigLIP2 image model achieved TSS=0.646, HSS=0.261, and MCC=0.340, confirming the ability of vision–language transformers to extract meaningful magnetic features from single magnetograms. The VideoMAE model, trained on 16-frame magnetogram sequences, yielded slightly lower scores (TSS=0.604, HSS=0.207, MCC=0.288), suggesting a limited ability to fully exploit temporal information and capture pre-flare evolution, even when varying cadence intervals (32, 72).
Overall, both the image and video models achieve slightly higher metric values when evaluated on the original data split compared to the alternative samples. The Moirai2 time-series model achieved the best overall performance, with TSS=0.736, HSS=0.691, and MCC=0.694, outperforming both the image and video models, particularly for the HSS and MCC metrics.
@@ Acknowledgements
The software code used in this work is publicly available under the GNU General Public License v3.0151515https://www.gnu.org/licenses/gpl-3.0.html on the GitHub repository https://github.com/inaf-oact-ai/solar-flare-forecaster/.
@@ Appendix B Forecasting metrics
where the recall (or true positive rate/sensitivity ) and the precision are respectively computed as:
@@
where is the true negative rate (specificity), computed as:
TSS ranges from (perfectly wrong) through (no skill, random/constant forecast) to (perfect). TSS is widely used in space-weather because it is relatively insensitive to class imbalance, while other metrics like the F1-score are more affected when the proportion of the minority class (rare event) changes.